OpenStudy (anonymous):

The angle of rotation of a regular polygon is 36 degrees. What is the name of the polygon? Regular Pentagon Regular Hexagon Regular Decagon Regular Octagon

OpenStudy (anonymous):

360 / 36 =?

OpenStudy (anonymous):

It looks like you have a formula here. \[\frac{360}{\textrm{number of sides}}=\textrm{angle of rotation}\] and also \[\frac{360}{\textrm{angle of rotation}}=\textrm{number of sides}\]

OpenStudy (anonymous):

Right.